1.a rectangular field 70m long and 50m wide has a path of uniform width around it . if the area of the path is 1024 sq.cm.find the width of the path
2.The base and height of a triangle are (x+3) and (2x-5)cm respectively if the area of the triangle is 20 sq.cm find x
3.The length of a rectangle is 5cm longer than its width and its area is 66 sq.cm find the perimeter of the rectangle

Answer by stanbon(75887)

1.a rectangular field 70m long and 50m wide has a path of uniform width around it . if the area of the path is 1024 sq.m.find the width of the path
---
Draw the picture of a rectangle inside a rectangle.
---
The area of the inner rectangle is 70*50 = 3500 sq. meters
---
Area of the outer rectangle = (70+2x)(50+2x) = 4X^2 + 240x + 3500
----
Equation:
Outer area - Inner area = 4x^2+ 240x = 1024
x^2+60x-256 = 0
(x-4)(x+64) = 0
---
Positive Solution:
x = 4 meters.
====================

2.The base and height of a triangle are (x+3) and (2x-5)cm respectively if the area of the triangle is 20 sq.cm find x
------------------
Area = (1/2)(base)(height)
20 cm^2 = (1/2)(x+3)(2x-5)
40 = 2x^2+x-15
---
2x^2+x-55 = 0
(x-5)(2x+11) = 0
Positive solution:
x = 5
-----
base = x+3 = 9 cm
height = 2x-5 = 5 cm
=====================
3.The length of a rectangle is 5cm longer than its width and its area is 66 sq.cm find the perimeter of the rectangle.
---
Let width = x
The length = x+5
---
Equation:
x(x+5) = 66
x^2+5x-66 = 0
(x+11)(x-6) = 0
---
Positive Solution:
x = 6 cm
---
Perimeter = 2(width + length)
P = 2(6 + 11)
P = 2(17)
Perimeter = 34 cm
======================
Cheers,
Stan H.
==============